Mars has an albedo of 15\%

15%. Given that it doesn't have much of an atmosphere, estimate the mean surface temperature on Mars. Note: the semimajor axis of the orbit of Mars is 1.52 AU.


1
Expert's answer
2021-09-24T09:25:40-0400

The illuminance generated by the Sun on the orbit of the Mars is

"E = \dfrac{L_{\odot}}{4\pi a^2}" .

The energy gathered by the martian surface every second is "(1-A)E\cdot \pi r^2" , where r is the radius of the Mars.

Due to the energy balance, the equal amount of energy is released every second. If we assume Mars to behave like a perfect black body, this energy is "4\pi r^2\sigma T^4" (Stephen-Boltzmann law).

"\dfrac{L_{\odot}}{4\pi a^2} \cdot(1- A)\pi r^2 = 4\pi r^2\sigma T^4,\\
\, \\
\dfrac{L_{\odot}}{4\pi a^2} \cdot(1- A)= 4\sigma T^4,\\
\,\\
T = \sqrt[4]{\dfrac{(1-A)L_{\odot}}{16\pi \sigma a^2}} ,\\
T = \sqrt[4]{\dfrac{(1-0.15)\cdot4\cdot10^{26}}{16\pi \cdot5.67\cdot10^{-8}\cdot(1.52\cdot1.5\cdot10^{11})^2}} = 219\,\mathrm{K}."